# A Survey of Protocol Fuzzing

### v2

@@ Acknowledgements.
This research is supported by the National Natural Science Foundation of China (Grant No. 62125205), the Nanyang Technological University (NTU)-DESAY SV Research Program under Grant 2018-0980, the National Research Foundation Singapore under its AI Singapore Programme (AISG2-RP-2020-019), the National Research Foundation through its National Satellite of Excellence in Trustworthy Software Systems (NSOE-TSS) project under the National Cybersecurity R&D (NCR) Grant award no. NRF2018NCR-NSOE003-0001.References
@@
- 802 (2020) 2020. IEEE Standard for Local and Metropolitan Area Networks–Port-Based Network Access Control. IEEE Std 802.1X-2020 (Revision of IEEE Std 802.1X-2010 Incorporating IEEE Std 802.1Xbx-2014 and IEEE Std 802.1Xck-2018) (2020), 1–289. https://doi.org/10.1109/IEEESTD.2020.9018454
- 802 (2021) 2021. IEEE Standard for Information Technology–Telecommunications and Information Exchange between Systems - Local and Metropolitan Area Networks–Specific Requirements - Part 11: Wireless LAN Medium Access Control (MAC) and Physical Layer (PHY) Specifications. IEEE Std 802.11-2020 (Revision of IEEE Std 802.11-2016) (2021), 1–4379. https://doi.org/10.1109/IEEESTD.2021.9363693
@@
- Aichernig et al. (2021) Bernhard K. Aichernig, Edi Muškardin, and Andrea Pferscher. 2021. Learning-Based Fuzzing of IoT Message Brokers. In 2021 14th IEEE Conference on Software Testing, Verification and Validation (ICST). 47–58. https://doi.org/10.1109/ICST49551.2021.00017
- Alliance (2015) ZigBee Alliance. 2015. ZigBee Specification. https://zigbeealliance.org/wp-content/uploads/2019/11/docs-05-3474-21-0csg-zigbee-specification.pdf
@@
- Alshmrany and Cordeiro (2020) Kaled M. Alshmrany and Lucas C. Cordeiro. 2020. Finding Security Vulnerabilities in Network Protocol Implementations. CoRR abs/2001.09592 (2020). arXiv:2001.09592 https://arxiv.org/abs/2001.09592
@@
- Ammann et al. (2024) Max Ammann, Lucca Hirschi, and Steve Kremer. 2024. DY Fuzzing: Formal Dolev-Yao Models Meet Cryptographic Protocol Fuzz Testing. In 45th IEEE Symposium on Security and Privacy.
@@
- Amusuo et al. (2023) Paschal C Amusuo, Ricardo Andrés Calvo Méndez, Zhongwei Xu, Aravind Machinery, and James C Davis. 2023. Systematically Detecting Packet Validation Vulnerabilities in Embedded Network Stacks. In 2023 38th IEEE/ACM International Conference on Automated Software Engineering (ASE). IEEE, 926–938.
@@
- Andronidis and Cadar (2022) Anastasios Andronidis and Cristian Cadar. 2022. SnapFuzz: An Efficient Fuzzing Framework for Network Applications. CoRR abs/2201.04048 (2022). arXiv:2201.04048 https://arxiv.org/abs/2201.04048
@@
- Androulaki et al. (2018) Elli Androulaki, Artem Barger, Vita Bortnikov, Christian Cachin, Konstantinos Christidis, Angelo De Caro, David Enyeart, Christopher Ferris, Gennady Laventman, Yacov Manevich, et al. 2018. Hyperledger fabric: a distributed operating system for permissioned blockchains. In Proceedings of the thirteenth EuroSys conference. 1–15.
@@
- Asadian et al. (2022) Hooman Asadian, Paul Fiterău-Broştean, Bengt Jonsson, and Konstantinos Sagonas. 2022. Applying Symbolic Execution to Test Implementations of a Network Protocol Against its Specification. In 2022 IEEE Conference on Software Testing, Verification and Validation (ICST). 70–81. https://doi.org/10.1109/ICST53961.2022.00019
@@
- Aschermann et al. (2020) Cornelius Aschermann, Sergej Schumilo, Ali Abbasi, and Thorsten Holz. 2020. Ijon: Exploring Deep State Spaces via Fuzzing. In 2020 IEEE Symposium on Security and Privacy (SP). 1597–1612. https://doi.org/10.1109/SP40000.2020.00117
@@
- Atlidakis et al. (2019) Vaggelis Atlidakis, Patrice Godefroid, and Marina Polishchuk. 2019. RESTler: stateful REST API fuzzing. In Proceedings of the 41st International Conference on Software Engineering, ICSE 2019, Montreal, QC, Canada, May 25-31, 2019, Joanne M. Atlee, Tevfik Bultan, and Jon Whittle (Eds.). IEEE / ACM, 748–758. https://doi.org/10.1109/ICSE.2019.00083
- AUTOSAR (2016) AUTOSAR. 2016. SOME/IP Protocol Specification. https://www.autosar.org/fileadmin/user_upload/standards/foundation/1-0/AUTOSAR_PRS_SOMEIPProtocol.pdf
@@
- Ba et al. (2022) Jinsheng Ba, Marcel Böhme, Zahra Mirzamomen, and Abhik Roychoudhury. 2022. Stateful Greybox Fuzzing. arXiv preprint arXiv:2204.02545 (2022).
@@
- Balakrishnan et al. (1995) Hari Balakrishnan, Srinivasan Seshan, Elan Amir, and Randy H Katz. 1995. Improving TCP/IP performance over wireless networks. In Proceedings of the 1st annual international conference on Mobile computing and networking. 2–11.
@@
- Bars et al. (2023) Nils Bars, Moritz Schloegel, Tobias Scharnowski, Nico Schiller, and Thorsten Holz. 2023. Fuzztruction: Using Fault Injection-based Fuzzing to Leverage Implicit Domain Knowledge. In USENIX Security.
@@
- Beurdouche et al. (2015) Benjamin Beurdouche, Karthikeyan Bhargavan, Antoine Delignat-Lavaud, Cédric Fournet, Markulf Kohlweiss, Alfredo Pironti, Pierre-Yves Strub, and Jean Karim Zinzindohoue. 2015. A Messy State of the Union: Taming the Composite State Machines of TLS. In 2015 IEEE Symposium on Security and Privacy. 535–552. https://doi.org/10.1109/SP.2015.39
@@
- Borcherding et al. (2023) Anne Borcherding, Mark Giraud, Ian Fitzgerald, and Jürgen Beyerer. 2023. The Bandit’s States: Modeling State Selection for Stateful Network Fuzzing as Multi-armed Bandit Problem. In 2023 IEEE European Symposium on Security and Privacy Workshops (EuroS&PW). IEEE, 345–350.
@@
- Bushart and Rossow (2023) Jonas Bushart and Christian Rossow. 2023. ResolFuzz: Differential Fuzzing of DNS Resolvers. ESORICS 2023 (2023).
@@
- Bytes et al. (2022) Andrei Bytes, Prashant Hari Narayan Rajput, Michail Maniatakos, and Jianying Zhou. 2022. FieldFuzz: Enabling vulnerability discovery in Industrial Control Systems supply chain using stateful system-level fuzzing. https://doi.org/10.48550/ARXIV.2204.13499
@@
- Castro et al. (1999) Miguel Castro, Barbara Liskov, et al. 1999. Practical byzantine fault tolerance. In OsDI, Vol. 99. 173–186.
- Chaki and Datta (2009) Sagar Chaki and Anupam Datta. 2009. ASPIER: An Automated Framework for Verifying Security Protocol Implementations. In 2009 22nd IEEE Computer Security Foundations Symposium. 172–185. https://doi.org/10.1109/CSF.2009.20
- Chau et al. (2017) Sze Yiu Chau, Omar Chowdhury, Endadul Hoque, Huangyi Ge, Aniket Kate, Cristina Nita-Rotaru, and Ninghui Li. 2017. SymCerts: Practical Symbolic Execution for Exposing Noncompliance in X.509 Certificate Validation Implementations. In 2017 IEEE Symposium on Security and Privacy (SP). 503–520. https://doi.org/10.1109/SP.2017.40
- Chen et al. (2018) Jiongyi Chen, Wenrui Diao, Qingchuan Zhao, Chaoshun Zuo, Zhiqiang Lin, XiaoFeng Wang, Wing Cheong Lau, Menghan Sun, Ronghai Yang, and Kehuan Zhang. 2018. IoTFuzzer: Discovering Memory Corruptions in IoT Through App-based Fuzzing. In 25th Annual Network and Distributed System Security Symposium, NDSS 2018, San Diego, California, USA, February 18-21, 2018. The Internet Society. http://wp.internetsociety.org/ndss/wp-content/uploads/sites/25/2018/02/ndss2018_01A-1_Chen_paper.pdf
@@
- Chen et al. (2021) Mark Chen, Jerry Tworek, Heewoo Jun, Qiming Yuan, Henrique Ponde de Oliveira Pinto, Jared Kaplan, Harri Edwards, Yuri Burda, Nicholas Joseph, Greg Brockman, et al. 2021. Evaluating large language models trained on code. arXiv preprint arXiv:2107.03374 (2021).
- Chen et al. (2019) Yurong Chen, Tian lan, and Guru Venkataramani. 2019. Exploring Effective Fuzzing Strategies to Analyze Communication Protocols. In Proceedings of the 3rd ACM Workshop on Forming an Ecosystem Around Software Transformation (London, United Kingdom) (FEAST’19). Association for Computing Machinery, New York, NY, USA, 17–23. https://doi.org/10.1145/3338502.3359762
- Chen et al. (2022) Yuanliang Chen, Fuchen Ma, Yuanhang Zhou, Yu Jiang, Ting Chen, and Jiaguang Sun. 2022. Tyr: Finding consensus failure bugs in blockchain system with behaviour divergent model. In 2023 IEEE Symposium on Security and Privacy (SP). IEEE Computer Society, 1186–1201.
- Chen et al. (2023) Yuanliang Chen, Fuchen Ma, Yuanhang Zhou, Yu Jiang, Ting Chen, and Jiaguang Sun. 2023. Tyr: Finding consensus failure bugs in blockchain system with behaviour divergent model. In 2023 IEEE Symposium on Security and Privacy (SP). IEEE, 2517–2532.
@@
- Comer (2013) Douglas E Comer. 2013. Internetworking with TCP/IP. Addison-Wesley Professional.
@@
- Dai et al. (2010) Huning Dai, Christian Murphy, and Gail Kaiser. 2010. Configuration Fuzzing for Software Vulnerability Detection. In 2010 International Conference on Availability, Reliability and Security. 525–530. https://doi.org/10.1109/ARES.2010.22
- Daniel et al. (2018) Lesly-Ann Daniel, Erik Poll, and Joeri de Ruiter. 2018. Inferring OpenVPN State Machines Using Protocol State Fuzzing. In 2018 IEEE European Symposium on Security and Privacy Workshops (EuroS&PW). 11–19. https://doi.org/10.1109/EuroSPW.2018.00009
- de Ruiter and Poll (2015) Joeri de Ruiter and Erik Poll. 2015. Protocol State Fuzzing of TLS Implementations. In 24th USENIX Security Symposium (USENIX Security 15). USENIX Association, Washington, D.C., 193–206. https://www.usenix.org/conference/usenixsecurity15/technical-sessions/presentation/de-ruiter
@@
- Du and Li (2023) Zhengjie Du and Yuekang Li. 2023. HasteFuzz: Full-Speed Fuzzing. In 2023 IEEE/ACM International Workshop on Search-Based and Fuzz Testing (SBFT). IEEE, 73–75.
@@
- Fan and Chang (2018) Rong Fan and Yaoyao Chang. 2018. Machine Learning for Black-Box Fuzzing of Network Protocols. In Information and Communications Security, Sihan Qing, Chris Mitchell, Liqun Chen, and Dongmei Liu (Eds.). Springer International Publishing, Cham, 621–632.
- Fang et al. (2021) Dongliang Fang, Zhanwei Song, Le Guan, Puzhuo Liu, Anni Peng, Kai Cheng, Yaowen Zheng, Peng Liu, Hongsong Zhu, and Limin Sun. 2021. ICS3Fuzzer: A Framework for Discovering Protocol Implementation Bugs in ICS Supervisory Software by Fuzzing. In ACSAC ’21: Annual Computer Security Applications Conference, Virtual Event, USA, December 6 - 10, 2021. ACM, 849–860. https://doi.org/10.1145/3485832.3488028
- Feng et al. (2021) Xiaotao Feng, Ruoxi Sun, Xiaogang Zhu, Minhui Xue, Sheng Wen, Dongxi Liu, Surya Nepal, and Yang Xiang. 2021. Snipuzz: Black-Box Fuzzing of IoT Firmware via Message Snippet Inference. In Proceedings of the 2021 ACM SIGSAC Conference on Computer and Communications Security (Virtual Event, Republic of Korea) (CCS ’21). Association for Computing Machinery, New York, NY, USA, 337–350. https://doi.org/10.1145/3460120.3484543
- Fiterau-Brostean et al. (2020) Paul Fiterau-Brostean, Bengt Jonsson, Robert Merget, Joeri de Ruiter, Konstantinos Sagonas, and Juraj Somorovsky. 2020. Analysis of DTLS Implementations Using Protocol State Fuzzing. In 29th USENIX Security Symposium (USENIX Security 20). USENIX Association, 2523–2540. https://www.usenix.org/conference/usenixsecurity20/presentation/fiterau-brostean
- Fiterau-Brostean et al. (2022) P. Fiterau-Brostean, B. Jonsson, K. Sagonas, and F. Taquist. 2022. DTLS-Fuzzer: A DTLS Protocol State Fuzzer. In 2022 IEEE Conference on Software Testing, Verification and Validation (ICST). IEEE Computer Society, Los Alamitos, CA, USA, 456–458. https://doi.org/10.1109/ICST53961.2022.00051
- Fiterau-Brostean et al. (2023) Paul Fiterau-Brostean, Bengt Jonsson, Konstantinos Sagonas, and Fredrik Tåquist. 2023. Automata-Based Automated Detection of State Machine Bugs in Protocol Implementations.. In NDSS.
@@
- Garbelini et al. (2022a) Matheus E. Garbelini, Vaibhav Bedi, Sudipta Chattopadhyay, Sumei Sun, and Ernest Kurniawan. 2022a. BrakTooth: Causing Havoc on Bluetooth Link Manager via Directed Fuzzing. In 31st USENIX Security Symposium (USENIX Security 22). USENIX Association, Boston, MA, 1025–1042. https://www.usenix.org/conference/usenixsecurity22/presentation/garbelini
- Garbelini et al. (2022b) Matheus E. Garbelini, Zewen Shang, Sudipta Chattopadhyay, Sumei Sun, and Ernest Kurniawan. 2022b. Towards Automated Fuzzing of 4G/5G Protocol Implementations Over the Air. In GLOBECOM 2022 - 2022 IEEE Global Communications Conference. 86–92. https://doi.org/10.1109/GLOBECOM48099.2022.10001673
- Garbelini et al. (2022c) Matheus E. Garbelini, Chundong Wang, and Sudipta Chattopadhyay. 2022c. Greyhound: Directed Greybox Wi-Fi Fuzzing. IEEE Transactions on Dependable and Secure Computing 19, 2 (2022), 817–834. https://doi.org/10.1109/TDSC.2020.3014624
- Garbelini et al. (2020) Matheus E. Garbelini, Chundong Wang, Sudipta Chattopadhyay, Sun Sumei, and Ernest Kurniawan. 2020. SweynTooth: Unleashing Mayhem over Bluetooth Low Energy. In 2020 USENIX Annual Technical Conference (USENIX ATC 20). USENIX Association, 911–925. https://www.usenix.org/conference/atc20/presentation/garbelini
- Gascon et al. (2015) Hugo Gascon, Christian Wressnegger, Fabian Yamaguchi, Daniel Arp, and Konrad Rieck. 2015. Pulsar: Stateful Black-Box Fuzzing of Proprietary Network Protocols. In Security and Privacy in Communication Networks - 11th International Conference, SecureComm 2015, Dallas, TX, USA, October 26-29, 2015, Revised Selected Papers (Lecture Notes of the Institute for Computer Sciences, Social Informatics and Telecommunications Engineering, Vol. 164), Bhavani M. Thuraisingham, XiaoFeng Wang, and Vinod Yegneswaran (Eds.). Springer, 330–347. https://doi.org/10.1007/978-3-319-28865-9_18
@@
- Goubault-Larrecq and Parrennes (2005) Jean Goubault-Larrecq and Fabrice Parrennes. 2005. Cryptographic Protocol Analysis on Real C Code. In Verification, Model Checking, and Abstract Interpretation, Radhia Cousot (Ed.). Springer Berlin Heidelberg, Berlin, Heidelberg, 363–379.
@@
- He et al. (2022) Fengjiao He, Wenchuan Yang, Baojiang Cui, and Jia Cui. 2022. Intelligent Fuzzing Algorithm for 5G NAS Protocol Based on Predefined Rules. In 2022 International Conference on Computer Communications and Networks (ICCCN). 1–7. https://doi.org/10.1109/ICCCN54977.2022.9868872
@@
- Heinze et al. (2020) Dennis Heinze, Jiska Classen, and Matthias Hollick. 2020. ToothPicker: Apple Picking in the iOS Bluetooth Stack. In 14th USENIX Workshop on Offensive Technologies (WOOT 20). USENIX Association. https://www.usenix.org/conference/woot20/presentation/heinze
@@
- Hoque et al. (2017) Endadul Hoque, Omar Chowdhury, Sze Yiu Chau, Cristina Nita-Rotaru, and Ninghui Li. 2017. Analyzing Operational Behavior of Stateful Protocol Implementations for Detecting Semantic Bugs. In 2017 47th Annual IEEE/IFIP International Conference on Dependable Systems and Networks (DSN). 627–638. https://doi.org/10.1109/DSN.2017.36
@@
- Hussain et al. (2021) Syed Rafiul Hussain, Imtiaz Karim, Abdullah Al Ishtiaq, Omar Chowdhury, and Elisa Bertino. 2021. Noncompliance as deviant behavior: An automated black-box noncompliance checker for 4g lte cellular devices. In Proceedings of the 2021 ACM SIGSAC Conference on Computer and Communications Security. 1082–1099.
@@
- Iyengar and Thomson (2021) Jana Iyengar and Martin Thomson. 2021. QUIC: A UDP-Based Multiplexed and Secure Transport. RFC 9000. https://doi.org/10.17487/RFC9000
@@
- Jelassi et al. (2012) Sofiene Jelassi, Gerardo Rubino, Hugh Melvin, Habib Youssef, and Guy Pujolle. 2012. Quality of Experience of VoIP Service: A Survey of Assessment Approaches and Open Issues. IEEE Communications Surveys & Tutorials 14, 2 (2012), 491–513. https://doi.org/10.1109/SURV.2011.120811.00063

### v3

@@ Acknowledgements.
This research is supported by the National Natural Science Foundation of China (Grant Nos. 62125205 and U23A20303), the ‘111 Center’ (B16037), the Key Research and Development Program of Shaanxi under Grant 2023KXJ190, and the Fundamental Research Funds for the Central Universities, No.YJSJ24010. The Nanyang Technological University (NTU)-DESAY SV Research Program under Grant 2018-0980, the National Research Foundation Singapore under its AI Singapore Programme (AISG2-RP-2020-019), the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN). Any opinions, findings, conclusions or recommendations expressed in this material are those of the authors and do not reflect the views of National Research Foundation, Singapore, and the Cyber Security Agency of Singapore.References
@@
- 802 (2020) 2020. IEEE Standard for Local and Metropolitan Area Networks–Port-Based Network Access Control. IEEE Std 802.1X-2020 (Revision of IEEE Std 802.1X-2010 Incorporating IEEE Std 802.1Xbx-2014 and IEEE Std 802.1Xck-2018) (2020), 1–289.
- 802 (2021) 2021. IEEE Standard for Information Technology–Telecommunications and Information Exchange between Systems - Local and Metropolitan Area Networks–Specific Requirements - Part 11: Wireless LAN Medium Access Control (MAC) and Physical Layer (PHY) Specifications. IEEE Std 802.11-2020 (Revision of IEEE Std 802.11-2016) (2021), 1–4379.
@@
- Aichernig et al. (2021) Bernhard K. Aichernig, Edi Muškardin, and Andrea Pferscher. 2021. Learning-Based Fuzzing of IoT Message Brokers. In 2021 14th IEEE Conference on Software Testing, Verification and Validation (ICST). 47–58.
- Alliance (2015) ZigBee Alliance. 2015. ZigBee Specification.
@@
- Alshmrany and Cordeiro (2020) Kaled M. Alshmrany and Lucas C. Cordeiro. 2020. Finding Security Vulnerabilities in Network Protocol Implementations. (2020). arXiv:2001.09592
@@
- Ammann et al. (2024) Max Ammann, Lucca Hirschi, and Steve Kremer. 2024. DY Fuzzing: Formal Dolev-Yao Models Meet Cryptographic Protocol Fuzz Testing. In 45th IEEE Symposium on Security and Privacy (SP). 99–99.
@@
- Amusuo et al. (2023) Paschal C Amusuo, Ricardo Andrés Calvo Méndez, Zhongwei Xu, Aravind Machinery, and James C Davis. 2023. Systematically Detecting Packet Validation Vulnerabilities in Embedded Network Stacks. In 2023 38th IEEE/ACM International Conference on Automated Software Engineering (ASE). 926–938.
@@
- Andronidis and Cadar (2022) Anastasios Andronidis and Cristian Cadar. 2022. SnapFuzz: An Efficient Fuzzing Framework for Network Applications. (2022). arXiv:2201.04048
@@
- Androulaki et al. (2018) Elli Androulaki, Artem Barger, Vita Bortnikov, Christian Cachin, Konstantinos Christidis, Angelo De Caro, David Enyeart, Christopher Ferris, Gennady Laventman, Yacov Manevich, et al. 2018. Hyperledger fabric: a distributed operating system for permissioned blockchains. In thirteenth EuroSys conference. 1–15.
@@
- Asadian et al. (2022) Hooman Asadian, Paul Fiterău-Broştean, Bengt Jonsson, and Konstantinos Sagonas. 2022. Applying Symbolic Execution to Test Implementations of a Network Protocol Against its Specification. In 2022 IEEE Conference on Software Testing, Verification and Validation (ICST). 70–81.
@@
- Aschermann et al. (2020) Cornelius Aschermann, Sergej Schumilo, Ali Abbasi, and Thorsten Holz. 2020. Ijon: Exploring Deep State Spaces via Fuzzing. In 2020 IEEE Symposium on Security and Privacy (SP). 1597–1612.
@@
- Atlidakis et al. (2019) Vaggelis Atlidakis, Patrice Godefroid, and Marina Polishchuk. 2019. RESTler: stateful REST API fuzzing. In 41st International Conference on Software Engineering (ICSE). 748–758.
- AUTOSAR (2016) AUTOSAR. 2016. SOME/IP Protocol Specification.
@@
- Ba et al. (2022) Jinsheng Ba, Marcel Böhme, Zahra Mirzamomen, and Abhik Roychoudhury. 2022. Stateful Greybox Fuzzing. In 31st USENIX Security Symposium (USENIX Security). 3255–3272.
@@
- Balakrishnan et al. (1995) Hari Balakrishnan, Srinivasan Seshan, Elan Amir, and Randy H Katz. 1995. Improving TCP/IP performance over wireless networks. In 1st annual international conference on Mobile computing and networking. 2–11.
@@
- Bars et al. (2023) Nils Bars, Moritz Schloegel, Tobias Scharnowski, Nico Schiller, and Thorsten Holz. 2023. Fuzztruction: Using Fault Injection-based Fuzzing to Leverage Implicit Domain Knowledge. In 32nd USENIX Security Symposium (USENIX Security 23). 1847–1864.
@@
- Beurdouche et al. (2015) Benjamin Beurdouche, Karthikeyan Bhargavan, Antoine Delignat-Lavaud, Cédric Fournet, Markulf Kohlweiss, Alfredo Pironti, Pierre-Yves Strub, and Jean Karim Zinzindohoue. 2015. A Messy State of the Union: Taming the Composite State Machines of TLS. In 2015 IEEE Symposium on Security and Privacy (SP). 535–552.
@@
- Borcherding et al. (2023) Anne Borcherding, Mark Giraud, Ian Fitzgerald, and Jürgen Beyerer. 2023. The Bandit’s States: Modeling State Selection for Stateful Network Fuzzing as Multi-armed Bandit Problem. In 2023 IEEE European Symposium on Security and Privacy Workshops (EuroS&PW). 345–350.
@@
- Bushart and Rossow (2023) Jonas Bushart and Christian Rossow. 2023. ResolFuzz: Differential Fuzzing of DNS Resolvers. In 28th European Symposium on Research in Computer Security (ESORICS). 62–80.
@@
- Bytes et al. (2022) Andrei Bytes, Prashant Hari Narayan Rajput, Michail Maniatakos, and Jianying Zhou. 2022. FieldFuzz: Enabling vulnerability discovery in Industrial Control Systems supply chain using stateful system-level fuzzing. arXiv:2204.13499
- Caballero et al. (2007) Juan Caballero, Heng Yin, Zhenkai Liang, and Dawn Song. 2007. Polyglot: automatic extraction of protocol message format using dynamic binary analysis. In 14th ACM Conference on Computer and Communications Security (CCS). 317–329.
@@
- Castro et al. (1999) Miguel Castro, Barbara Liskov, et al. 1999. Practical byzantine fault tolerance. In 3rd Symposium on Operating Systems Design and Implementation (OSDI), Vol. 99. 173–186.
- Chaki and Datta (2009) Sagar Chaki and Anupam Datta. 2009. ASPIER: An Automated Framework for Verifying Security Protocol Implementations. In 2009 22nd IEEE Computer Security Foundations Symposium. 172–185.
- Chau et al. (2017) Sze Yiu Chau, Omar Chowdhury, Endadul Hoque, Huangyi Ge, Aniket Kate, Cristina Nita-Rotaru, and Ninghui Li. 2017. SymCerts: Practical Symbolic Execution for Exposing Noncompliance in X.509 Certificate Validation Implementations. In 2017 IEEE Symposium on Security and Privacy (SP). 503–520.
- Chen et al. (2018) Jiongyi Chen, Wenrui Diao, Qingchuan Zhao, Chaoshun Zuo, Zhiqiang Lin, XiaoFeng Wang, Wing Cheong Lau, Menghan Sun, Ronghai Yang, and Kehuan Zhang. 2018. IoTFuzzer: Discovering Memory Corruptions in IoT Through App-based Fuzzing. In 25th Annual Network and Distributed System Security Symposium (NDSS).
@@
- Chen et al. (2019) Yurong Chen, Tian lan, and Guru Venkataramani. 2019. Exploring Effective Fuzzing Strategies to Analyze Communication Protocols. In 3rd ACM Workshop on Forming an Ecosystem Around Software Transformation (FEAST). 17–23.
- Chen et al. (2023) Yuanliang Chen, Fuchen Ma, Yuanhang Zhou, Yu Jiang, Ting Chen, and Jiaguang Sun. 2023. Tyr: Finding Consensus Failure Bugs in Blockchain System with Behaviour Divergent Model. In 2023 IEEE Symposium on Security and Privacy (SP). 2517–2532.
@@
- Comer (2013) Douglas E. Comer. 2013. Internetworking with TCP/IP (6th ed.). Addison-Wesley Professional.
- Comparetti et al. (2009) Paolo Milani Comparetti, Gilbert Wondracek, Christopher Kruegel, and Engin Kirda. 2009. Prospex: Protocol Specification Extraction. In 2009 30th IEEE Symposium on Security and Privacy. 110–125.
@@
- Dai et al. (2010) Huning Dai, Christian Murphy, and Gail Kaiser. 2010. Configuration Fuzzing for Software Vulnerability Detection. In 2010 International Conference on Availability, Reliability and Security. 525–530.
- Daniel et al. (2018) Lesly-Ann Daniel, Erik Poll, and Joeri de Ruiter. 2018. Inferring OpenVPN State Machines Using Protocol State Fuzzing. In 2018 IEEE European Symposium on Security and Privacy Workshops (EuroS&PW). 11–19.
- de Ruiter and Poll (2015) Joeri de Ruiter and Erik Poll. 2015. Protocol State Fuzzing of TLS Implementations. In 24th USENIX Security Symposium (USENIX Security). 193–206.
@@
- Fan and Chang (2018) Rong Fan and Yaoyao Chang. 2018. Machine Learning for Black-Box Fuzzing of Network Protocols. In Information and Communications Security. 621–632.
- Fang et al. (2021) Dongliang Fang, Zhanwei Song, Le Guan, Puzhuo Liu, Anni Peng, Kai Cheng, Yaowen Zheng, Peng Liu, Hongsong Zhu, and Limin Sun. 2021. ICS3Fuzzer: A Framework for Discovering Protocol Implementation Bugs in ICS Supervisory Software by Fuzzing. In Annual Computer Security Applications Conference (ACSAC). 849–860.
- Feng et al. (2021) Xiaotao Feng, Ruoxi Sun, Xiaogang Zhu, Minhui Xue, Sheng Wen, Dongxi Liu, Surya Nepal, and Yang Xiang. 2021. Snipuzz: Black-Box Fuzzing of IoT Firmware via Message Snippet Inference. In 2021 ACM SIGSAC Conference on Computer and Communications Security (CCS). 337–350.
- Fiterau-Brostean et al. (2020) Paul Fiterau-Brostean, Bengt Jonsson, Robert Merget, Joeri de Ruiter, Konstantinos Sagonas, and Juraj Somorovsky. 2020. Analysis of DTLS Implementations Using Protocol State Fuzzing. In 29th USENIX Security Symposium (USENIX Security). 2523–2540.
- Fiterau-Brostean et al. (2022) P. Fiterau-Brostean, B. Jonsson, K. Sagonas, and F. Taquist. 2022. DTLS-Fuzzer: A DTLS Protocol State Fuzzer. In 2022 IEEE Conference on Software Testing, Verification and Validation (ICST). 456–458.
- Fiterau-Brostean et al. (2023) Paul Fiterau-Brostean, Bengt Jonsson, Konstantinos Sagonas, and Fredrik Tåquist. 2023. Automata-Based Automated Detection of State Machine Bugs in Protocol Implementations.. In 30rd Annual Network and Distributed System Security Symposium (NDSS).
- Foundation (2015) Open Networking Foundation. 2015. OpenFlow Switch Specification.
@@
- Garbelini et al. (2022a) Matheus E. Garbelini, Vaibhav Bedi, Sudipta Chattopadhyay, Sumei Sun, and Ernest Kurniawan. 2022a. BrakTooth: Causing Havoc on Bluetooth Link Manager via Directed Fuzzing. In 31st USENIX Security Symposium (USENIX Security 22). 1025–1042.
- Garbelini et al. (2022b) Matheus E. Garbelini, Zewen Shang, Sudipta Chattopadhyay, Sumei Sun, and Ernest Kurniawan. 2022b. Towards Automated Fuzzing of 4G/5G Protocol Implementations Over the Air. In GLOBECOM 2022 - 2022 IEEE Global Communications Conference. 86–92.
- Garbelini et al. (2022c) Matheus E. Garbelini, Chundong Wang, and Sudipta Chattopadhyay. 2022c. Greyhound: Directed Greybox Wi-Fi Fuzzing. IEEE Transactions on Dependable and Secure Computing 19, 2 (2022), 817–834.
- Garbelini et al. (2020) Matheus E. Garbelini, Chundong Wang, Sudipta Chattopadhyay, Sun Sumei, and Ernest Kurniawan. 2020. SweynTooth: Unleashing Mayhem over Bluetooth Low Energy. In 2020 USENIX Annual Technical Conference (USENIX ATC 20). 911–925.
- Gascon et al. (2015) Hugo Gascon, Christian Wressnegger, Fabian Yamaguchi, Daniel Arp, and Konrad Rieck. 2015. Pulsar: Stateful Black-Box Fuzzing of Proprietary Network Protocols. In Security and Privacy in Communication Networks - 11th International Conference (SecureComm), Vol. 164. 330–347.
@@
- Goubault-Larrecq and Parrennes (2005) Jean Goubault-Larrecq and Fabrice Parrennes. 2005. Cryptographic Protocol Analysis on Real C Code. In Verification, Model Checking, and Abstract Interpretation. 363–379.
@@
- He et al. (2022) Fengjiao He, Wenchuan Yang, Baojiang Cui, and Jia Cui. 2022. Intelligent Fuzzing Algorithm for 5G NAS Protocol Based on Predefined Rules. In 2022 International Conference on Computer Communications and Networks (ICCCN). 1–7.
@@
- Heinze et al. ([n. d.]) Dennis Heinze, Jiska Classen, and Matthias Hollick. [n. d.]. ToothPicker: Apple Picking in the iOS Bluetooth Stack. In 14th USENIX Workshop on Offensive Technologies (WOOT).
@@
- Hoque et al. (2017) Endadul Hoque, Omar Chowdhury, Sze Yiu Chau, Cristina Nita-Rotaru, and Ninghui Li. 2017. Analyzing Operational Behavior of Stateful Protocol Implementations for Detecting Semantic Bugs. In 47th IEEE/IFIP International Conference on Dependable Systems and Networks (DSN). 627–638.
@@
- Hussain et al. (2021) Syed Rafiul Hussain, Imtiaz Karim, Abdullah Al Ishtiaq, Omar Chowdhury, and Elisa Bertino. 2021. Noncompliance as deviant behavior: An automated black-box noncompliance checker for 4g lte cellular devices. In 2021 ACM SIGSAC Conference on Computer and Communications Security (CCS). 1082–1099.
@@
- Iyengar and Thomson (2021) Jana Iyengar and Martin Thomson. 2021. QUIC: A UDP-Based Multiplexed and Secure Transport. RFC 9000. https://www.rfc-editor.org/info/rfc9000
@@
- Jelassi et al. (2012) Sofiene Jelassi, Gerardo Rubino, Hugh Melvin, Habib Youssef, and Guy Pujolle. 2012. Quality of Experience of VoIP Service: A Survey of Assessment Approaches and Open Issues. IEEE Communications Surveys & Tutorials 14, 2 (2012), 491–513.